OpenStudy (anonymous):

Choose the correct simplification of the expression the quantity ( I will write the quantity in a minute)

OpenStudy (anonymous):

\[\left(\begin{matrix}4x \\ y\end{matrix}\right)^{2}\]

OpenStudy (anonymous):

A. \[\frac{ 4x^{2} }{ y^{2} }\] B \[\frac{ 8x^{2} }{ y^{2} }\] C. \[\frac{ 16x^{2} }{ y^{2} }\] D. \[\frac{ 16x^{2} }{ y }\]

OpenStudy (anonymous):

C

OpenStudy (anonymous):

How is it C. Can I have a explination

OpenStudy (anonymous):

And can you help me with another?

OpenStudy (anonymous):

\[{\left( \frac{ 4x }{ y } \right)}^{2} = \frac{ 4²x² }{ y² } = \frac{ 16x² }{ y² }\]

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Ok here it is

OpenStudy (anonymous):

Choose the correct simplification of the expression \[\frac{ a^{6}b^{3} }{ a^{4}b^{2} }\]

OpenStudy (anonymous):

A. \[\frac{ a^{2} }{ b }\] B. \[a^{10}b^{5}\] C. \[\frac{ a^{10} }{ b^{5}}\] D. \[a^{2}b\]

OpenStudy (anonymous):

\[\frac{ a^6 b^3 }{ a^4 b^2 } =a^2 b\]

OpenStudy (anonymous):

explanation:

OpenStudy (anonymous):

Choose the correct simplification of the expression \[\frac{ 9 }{ h^{-3} }\]

OpenStudy (anonymous):

this one?

OpenStudy (anonymous):

\[\frac{ a \times a \times a \times a \times a \times a \times b \times b \times b}{ a \times a \times a \times a \times b \times b} = a^2 b\]

OpenStudy (anonymous):

\[\frac{ 9 }{ h^-2 } = 9h^2\]

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (anonymous):

wc ^^

OpenStudy (anonymous):

I have 1 more left

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Show the expanded form and simplification of \[\frac{ x^{6} }{ x^{2} }\] Explain in your own words how you can simplify \[\frac{ x^{6} }{ x^{2} }\] Create your own fraction with like bases, coefficients, and show its simplification.

OpenStudy (anonymous):

\[\frac{ x \times x \times x \times x \times x \times x }{ x \times x } = x^4\]

OpenStudy (anonymous):

\[\frac{ x^6 }{ x^2 } = x^{6-2} = x^4\]

OpenStudy (anonymous):

In general, \[\frac{ x^a }{ x^b } = x^{a-b}\]

OpenStudy (anonymous):

for a better understanding of all these operations i recommend this reading: https://en.wikipedia.org/wiki/Exponentiation

OpenStudy (anonymous):

thanks so much you have a great day

OpenStudy (anonymous):

thks, u too